OpenStudy (anonymous):

solve using the multiplication principle. -5/6x= -9/10 What is the solution? x=?

OpenStudy (anonymous):

can somebody help me??

OpenStudy (anonymous):

you need to take x off the denominator. so multiply everything by 6x. Then, you need to also destroy the other denominator, so multiply everything by 10. If you like, do a single multiplication with 60x

OpenStudy (anonymous):

I dont understand

OpenStudy (anonymous):

grr.. basics. ok. -5/6x=-9/10, 6x(-5/6x)=6x(-9/10), -5=-54x/10, -50=10(-54x/10), -50=-540x, 50=540x, x=5/54

OpenStudy (anonymous):

dont get angry please. I have had to learn nine things in a nine week course. I am not a math taker.